# The Promise and Challenges of Using LLMs to Accelerate the Screening Process of Systematic Reviews

### v2

@@ Abstract.
journalyear: 2024conference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italycopyright: none
@@ 1. Introduction
Systematic review (SR) is a popular research method in software engineering; since 1994, over a hundred SRs about software testing (garousiSystematicLiteratureReview2016, ) have been published. By a recent estimate, conducting a SR takes about 67 weeks (marshallSystematicReviewAutomation2019, ). A significant amount of time is spent on manual labor, which prior studies have tried to automate  (vandinterAutomationSystematicLiterature2021, ; marshallSystematicReviewAutomation2019, ; guoAutomatedPaperScreening2024, ; khraishaCanLargeLanguage2024, ; bolanosArtificialIntelligenceLiterature2024, ; wilkinsAutomatedTitleAbstract2023, ; robinsonBioSIEVEExploringInstruction2023, ; wangZeroshotGenerativeLarge2024, ). Our motivation for this study is to investigate whether the effort required for SRs can be reduced by using Large Language Models (LLMs) in text simplification or automating the whole screening process.
The SR process defines a screening process by which primary studies are included or excluded (kitchenhamGuidelinesPerformingSystematic2007, ). This process usually has two phases, the first using just the title and abstract; and the second using the full text. Title-abstract screening is usually faster and done before full-text screening (kitchenhamGuidelinesPerformingSystematic2007, ). The ability of LLMs to do complex language tasks makes them good candidates for speeding up the screening based on the titles and abstracts (zhaoSurveyLargeLanguage2023, ). Similarly, the ability of LLMs to do text simplification, which has been shown to increase reading comprehension  (crossley2014s, ), suggests their potential for improving screening performance and reducing manual labor in the screening process. We, therefore, pose the following research question:
@@ 2.1. Systematic Review Procedures
SRs are secondary research. Primary studies are those in which the authors generate or collect data. Secondary research aggregates the findings of primary studies (ralph2022paving, ). Many types of SRs exist, and protocols vary by type; however, SR protocols typically define:
@@
    A screening process by which selection criteria are applied to classify primary studies as included or excluded (kitchenhamGuidelinesPerformingSystematic2007, ; ralph2022paving, )
@@
Overall, conducting SRs can be time- and labor-intensive (vandinterAutomationSystematicLiterature2021, ; marshallSystematicReviewAutomation2019, ). Any technique that can exclude papers with high accuracy may therefore significantly accelerate the screening process.
@@ 2.2. Text Simplification
Text simplification is used to improve the readability and understandability of text while retaining its original meaning (al2021automated, ). Text simplification is often used with second language learners, children, people with disabilities, and people with low literacy (vstajner2021automatic, ). It seems plausible that text simplification could improve abstract comprehension by human screeners, especially for non-native English speakers.
@@ 2.3. Scientific Literacy
The ability to correctly screen primary studies depends on the screener’s scientific literacy. Scientific literacy is the ability to “use evidence and data to evaluate the quality of science information and arguments put forth by scientists and in the media” (gormally2012developing, ). Put another way, it is “the capacity to use scientific knowledge to identify questions and to draw evidence-based conclusions in order to understand and help make decisions about the natural world and the changes made to it through human activity” (gormally2012developing, ). Scientific literacy skills thus concern the use of data, evaluating the quality of scientific literature, making decisions, and reaching conclusions. Scientific literacy can be evaluated using a standardized test such as the Test Of Scientific Literacy Skills (TOSLS) (gormally2012developing, ).
@@ 2.4. Using Software to Help Systematic Reviews
Software should be able to automate or accelerate several aspects of SRs including constructing search strings, retrieving PDFs, extracting data, and automating screening. However, most existing studies focus on automating primary study screening and acknowledge that the main goal of automation is to reduce manual labor (vandinterAutomationSystematicLiterature2021, ). Prior work on using software to facilitate SRs focuses on reducing manual effort but stops short of automating the whole review process. For example, a review of text mining techniques used on SRs found that text mining techniques could reduce effort by 30–70%, but led to a $\approx 5\%$ loss of relevant studies (o2015using, ).
Marshall and Wallace (marshallSystematicReviewAutomation2019, ) provide a practical guide for using machine learning (ML) tools for literature search, screening, data extraction, or bias assessment. Several abstract-screening tools are available, of which Abstrackr and Research Screener tools reduce screening workload by 9–89%, with precision ranging from 16% to 45% (chaiResearchScreenerMachine2021, ; gatesTechnologyassistedTitleAbstract2018, ; rathboneFasterTitleAbstract2015a, ). All of these tools require a human in the loop, instructing the ML tool to identify relevant research papers.
Recent studies on using LLMs to automate the screening process in SRs have shown that LLMs can match human performance in paper screening, with some limitations (guoAutomatedPaperScreening2024, ; khraishaCanLargeLanguage2024, ; bolanosArtificialIntelligenceLiterature2024, ; wilkinsAutomatedTitleAbstract2023, ; robinsonBioSIEVEExploringInstruction2023, ; wangZeroshotGenerativeLarge2024, ). However, most of these approaches have been studied in the domain of medical sciences, the findings suggest that research on automating title-abstract screening could also be applied to other domains, such as SE.
@@ 2.5. Prompt Optimization
Incorporating task-specific examples and specially crafted instructions into LLM prompts has been shown to improve their performance in specific tasks (weiChainofThoughtPromptingElicits2023, ; kojimaLargeLanguageModels2023, ). Zero-shot (ZS) prompting refers to situations where no task-specific examples are provided. One-shot (OS) and Few-shot (FS) prompting are typical prompting techniques in natural language processing (NLP), providing task-specific examples (kojimaLargeLanguageModels2023, ). OS prompting incorporates one task example, while FS prompting includes multiple examples, with the number of examples depending on the context size of the LLM (brownLanguageModelsAre2020, ). Chain-of-Thought (CoT) prompting incorporates a “Let’s think step by step” phrase at the end of the prompt, instructing the LLM to decompose the task into smaller sub-tasks for more accurate task solving (weiChainofThoughtPromptingElicits2023, ). CoT prompting can be combined with FS prompting, which is known as Few-shot with Chain-of-Thought (FS-CoT) prompting, to further improve task accuracy (weiChainofThoughtPromptingElicits2023, ).
@@ 3.2. Abstract Simplification
From a pool of 5,414 papers originally screened for a prior SR (kuutila_time_2020, ), we selected 10 papers that were eventually included and 10 that were eventually excluded (top-left of Figure 1). We used our expertise to purposefully select difficult-to-screen papers. All selected papers have time-related terminology in their abstracts and mention empirical results related to time pressure. In the excluded papers, time pressure is used as a motivating factor for building something (e.g., testing framework), or time-related terminology is used in relation to other concepts such as temporal dispersion.
The inclusion criteria in the prior SR were:
(1) “The main focus is time pressure in software engineering“; and (2) “The paper presents empirical evidence of time pressure in software engineering“. (kuutila_time_2020, )
@@
Table 1 presents the lengths of the original and simplified abstracts and two readability metrics. The readability metrics were calculated using version 3.3.1 of the Quanteda library for the R programming language (benoit2018, ). The Flesch-Kincaid Grade level (FK) (kincaid1975derivation, ) and Simple Measure of Gobbledygook (SMOG) (mc1969smog, ) are readability metrics that estimate the US educational grade level (FK) or years of education (SMOG) required to completely understand a text. That is, lower scores indicate easier-to-read text. All selected papers and their original and simplified abstracts can be found in the replication package (see Data Availability).
@@ 3.4. LLM Reproduction of Title-Abstract Screening
We reproduced the experiment from Section 3.3 using LLMs, as shown in Figure 1. We used the gpt-3.5-turbo-0613 and gpt-4-0613 chat model implementations, as GPT-3.5 and GPT-4 are commonly used LLMs in SE-related research (houLargeLanguageModels2023, ). The corresponding LLM configurations are shown in Table 3. As the prompts in this experiment fit in the context of both models, the 16K context version (gpt-3.5-turbo-16k-0613) is unnecessary. We used a temperature value of zero to minimize the non-deterministic behavior of the tested LLMs, although some variation remains visible. The model cut-off date for all LLMs was set to June 2023 and a seed parameter of 128 was used in this and consequent experiments to make comparing the results more coherent and reproducible.
The reproduction uses the same screening task set of 20 papers from the prior SR  (kuutila_time_2020, ), as in the experiment done with humans. We constructed a prompt template following the screening procedure of the prior SR as closely as possible. Using this template, we generated prompts using the ZS prompting technique (see Section 2.5) to use in the reproduction. These prompts were then processed with a Python script, which called the OpenAI API. The prompt template, the generated prompts, the Python script, and the API responses are provided in the replication package (see Data Availability).
@@ 3.5. LLM Reproduction With Prompt Optimization
We selected 20 additional papers from the prior SR (kuutila_time_2020, ) as “seed papers”, of which 10 should be included and 10 should be excluded. The seed papers were used as examples with OS, FS, and FS-CoT prompting (see Section 2.5).
@@
For FS-CoT prompting, we reused the FS prompts. The phrase “Let’s think step by step” was appended to the end of each prompt. This phrase is used in Chain-of-Thought (CoT) prompting and improves task accuracy (weiChainofThoughtPromptingElicits2023, ; kojimaLargeLanguageModels2023, ).
@@ 3.8. Reproducing a Larger Screening Procedure
We also conducted exploratory analysis on 1,361 Scopus Papers that were initially gathered for the prior SR  (kuutila_time_2020, ) to see how LLMs perform with a more realistic sample size (see Figure 1). The Scopus Papers do not encompass all the screened articles for the prior SR, as the screened articles also included those from Google Scholar and articles resulting from snowballing, but it is the data that had complete abstracts readily available for analysis. Articles without abstracts or to which we did not have access were removed, leaving us with 1,306 papers.
@@ 4.1.2. Exploratory Analysis
Table 7. Human screener majority vote per paper. Paper references are given in the replication package  (zenodo, ).
@@
Judgment in
Original (kuutila_time_2020, )
Exclude Include
@@ 4.2. LLM Screening
Correct (%)
Model
Exclude Include Precision Recall
@@
Reference:
Human performance
@@ 5. Discussion
Paid researchers outperformed students in the screening task. This result is partly in conflict with the results of the experiment conducted by Salman et al. (salman2015students, ), where it was concluded that student performance was similar to professional performance in a programming task when the approach used was novel for both groups. In our case, the students had been taught about conducting secondary studies during the course they were attending, thus it was not completely novel for them, we did not record the experience of the researchers in performing them. Nevertheless, a noticeable result from the experiment with human screeners is that being a researcher and having higher scientific literacy skills, measured with the TOSLS, has a predictive ability on the number of correct judgments according to our regression model.
@@
The two existing ML tools, Abstrackr and Research Screener (covered in Section 2.4) had a screening accuracy between 16% and 45% (chaiResearchScreenerMachine2021, ; gatesTechnologyassistedTitleAbstract2018, ; rathboneFasterTitleAbstract2015a, ). Compared to our exploratory testing on a larger screening procedure, in Table 9, our screening performance seems significantly higher than with Abstrackr and Research Screener. However, the results are not directly comparable, as the datasets used were different, and Abstrackr gives recommendations instead of judgments. Also of note is that we picked 20 screening tasks for our experiments, and the performance for these particularly difficult cases is between 50% and 75% with a recall between 0.3 to 0.9. Precision and recall for each LLM and prompt combination is given in the replication package (zenodo, ). The recent studies on LLMs for automating the screening process of SRs (guoAutomatedPaperScreening2024, ; khraishaCanLargeLanguage2024, ; bolanosArtificialIntelligenceLiterature2024, ; wilkinsAutomatedTitleAbstract2023, ; robinsonBioSIEVEExploringInstruction2023, ; wangZeroshotGenerativeLarge2024, ) show that the accuracy of title-abstract screening with LLMs have varied between 0.60 and 0.835, which are similar to our results. Similar to the studies done with Abstrackr and Research Screener, our results are not directly comparable with the recent studies on LLMs, because of the datasets that focus on different aspects. Additionally, not all studies mention the used LLM versions, seed or temperature parameters making comparisons harder.
@@ 6. Limitations
As conducting SRs includes more than just paper screening, another avenue for future work is considering the use of LLMs in more phases. This includes the identification of studies for the screening (wohlin2022successful, ), data extraction and analysis.
@@ 7. Conclusions
The datasets and source code for the experiments are available at Zenodo  (zenodo, ). However, the data involving human screeners is not shared due to the absence of explicit consent for distribution.
Acknowledgements

### v3

@@ Abstract.
copyright: noneconference: 28th International Conference on Evaluation and Assessment in Software Engineering; June 18–21, 2024; Salerno, Italybooktitle: 28th International Conference on Evaluation and Assessment in Software Engineering (EASE 2024), June 18–21, 2024, Salerno, Italycopyright: none
@@ 1. Introduction
Systematic review (SR) is a popular research method in software engineering; since 1994, over a hundred SRs about software testing (Garousi and Mäntylä, 2016) have been published. By a recent estimate, conducting a SR takes about 67 weeks (Marshall and Wallace, 2019). A significant amount of time is spent on manual labor, which prior studies have tried to automate  (van Dinter et al., 2021; Marshall and Wallace, 2019; Guo et al., 2024; Khraisha et al., 2024; Bolanos et al., 2024; Wilkins, 2023; Robinson et al., 2023; Wang et al., 2024). Our motivation for this study is to investigate whether the effort required for SRs can be reduced by using Large Language Models (LLMs) in text simplification or automating the whole screening process.
The SR process defines a screening process by which primary studies are included or excluded (Kitchenham and Charters, 2007). This process usually has two phases, the first using just the title and abstract; and the second using the full text. Title-abstract screening is usually faster and done before full-text screening (Kitchenham and Charters, 2007). The ability of LLMs to do complex language tasks makes them good candidates for speeding up the screening based on the titles and abstracts (Zhao et al., 2023). Similarly, the ability of LLMs to do text simplification, which has been shown to increase reading comprehension  (Crossley et al., 2014), suggests their potential for improving screening performance and reducing manual labor in the screening process. We, therefore, pose the following research question:
@@ 2.1. Systematic Review Procedures
SRs are secondary research. Primary studies are those in which the authors generate or collect data. Secondary research aggregates the findings of primary studies (Ralph and Baltes, 2022). Many types of SRs exist, and protocols vary by type; however, SR protocols typically define:
@@
    A screening process by which selection criteria are applied to classify primary studies as included or excluded (Kitchenham and Charters, 2007; Ralph and Baltes, 2022)
@@
Overall, conducting SRs can be time- and labor-intensive (van Dinter et al., 2021; Marshall and Wallace, 2019). Any technique that can exclude papers with high accuracy may therefore significantly accelerate the screening process.
@@ 2.2. Text Simplification
Text simplification is used to improve the readability and understandability of text while retaining its original meaning (Al-Thanyyan and Azmi, 2021). Text simplification is often used with second language learners, children, people with disabilities, and people with low literacy (Štajner, 2021). It seems plausible that text simplification could improve abstract comprehension by human screeners, especially for non-native English speakers.
@@ 2.3. Scientific Literacy
The ability to correctly screen primary studies depends on the screener’s scientific literacy. Scientific literacy is the ability to “use evidence and data to evaluate the quality of science information and arguments put forth by scientists and in the media” (Gormally et al., 2012). Put another way, it is “the capacity to use scientific knowledge to identify questions and to draw evidence-based conclusions in order to understand and help make decisions about the natural world and the changes made to it through human activity” (Gormally et al., 2012). Scientific literacy skills thus concern the use of data, evaluating the quality of scientific literature, making decisions, and reaching conclusions. Scientific literacy can be evaluated using a standardized test such as the Test Of Scientific Literacy Skills (TOSLS) (Gormally et al., 2012).
@@ 2.4. Using Software to Help Systematic Reviews
Software should be able to automate or accelerate several aspects of SRs including constructing search strings, retrieving PDFs, extracting data, and automating screening. However, most existing studies focus on automating primary study screening and acknowledge that the main goal of automation is to reduce manual labor (van Dinter et al., 2021). Prior work on using software to facilitate SRs focuses on reducing manual effort but stops short of automating the whole review process. For example, a review of text mining techniques used on SRs found that text mining techniques could reduce effort by 30–70%, but led to a $\approx 5\%$ loss of relevant studies (O’Mara-Eves et al., 2015).
Marshall and Wallace (Marshall and Wallace, 2019) provide a practical guide for using machine learning (ML) tools for literature search, screening, data extraction, or bias assessment. Several abstract-screening tools are available, of which Abstrackr and Research Screener tools reduce screening workload by 9–89%, with precision ranging from 16% to 45% (Chai et al., 2021; Gates et al., 2018; Rathbone et al., 2015). All of these tools require a human in the loop, instructing the ML tool to identify relevant research papers.
Recent studies on using LLMs to automate the screening process in SRs have shown that LLMs can match human performance in paper screening, with some limitations (Guo et al., 2024; Khraisha et al., 2024; Bolanos et al., 2024; Wilkins, 2023; Robinson et al., 2023; Wang et al., 2024). However, most of these approaches have been studied in the domain of medical sciences, the findings suggest that research on automating title-abstract screening could also be applied to other domains, such as SE.
@@ 2.5. Prompt Optimization
Incorporating task-specific examples and specially crafted instructions into LLM prompts has been shown to improve their performance in specific tasks (Wei et al., 2023; Kojima et al., 2022). Zero-shot (ZS) prompting refers to situations where no task-specific examples are provided. One-shot (OS) and Few-shot (FS) prompting are typical prompting techniques in natural language processing (NLP), providing task-specific examples (Kojima et al., 2022). OS prompting incorporates one task example, while FS prompting includes multiple examples, with the number of examples depending on the context size of the LLM (Brown et al., 2020). Chain-of-Thought (CoT) prompting incorporates a “Let’s think step by step” phrase at the end of the prompt, instructing the LLM to decompose the task into smaller sub-tasks for more accurate task solving (Wei et al., 2023). CoT prompting can be combined with FS prompting, which is known as Few-shot with Chain-of-Thought (FS-CoT) prompting, to further improve task accuracy (Wei et al., 2023).
@@ 3.2. Abstract Simplification
From a pool of 5,414 papers originally screened for a prior SR (Kuutila et al., 2020), we selected 10 papers that were eventually included and 10 that were eventually excluded (top-left of Figure 1). We used our expertise to purposefully select difficult-to-screen papers. All selected papers have time-related terminology in their abstracts and mention empirical results related to time pressure. In the excluded papers, time pressure is used as a motivating factor for building something (e.g., testing framework), or time-related terminology is used in relation to other concepts such as temporal dispersion.
The inclusion criteria in the prior SR were: (1) “The main focus is time pressure in software engineering“; and (2) “The paper presents empirical evidence of time pressure in software engineering“. (Kuutila et al., 2020)
@@
Table 1 presents the lengths of the original and simplified abstracts and two readability metrics. The readability metrics were calculated using version 3.3.1 of the Quanteda library for the R programming language (Benoit et al., 2018). The Flesch-Kincaid Grade level (FK) (Kincaid et al., 1975) and Simple Measure of Gobbledygook (SMOG) (Mc Laughlin, 1969) are readability metrics that estimate the US educational grade level (FK) or years of education (SMOG) required to completely understand a text. That is, lower scores indicate easier-to-read text. All selected papers and their original and simplified abstracts can be found in the replication package (see Data Availability).
@@ 3.4. LLM Reproduction of Title-Abstract Screening
We reproduced the experiment from Section 3.3 using LLMs, as shown in Figure 1. We used the gpt-3.5-turbo-0613 and gpt-4-0613 chat model implementations, as GPT-3.5 and GPT-4 are commonly used LLMs in SE-related research (Hou et al., 2023). The corresponding LLM configurations are shown in Table 3. As the prompts in this experiment fit in the context of both models, the 16K context version (gpt-3.5-turbo-16k-0613) is unnecessary. We used a temperature value of zero to minimize the non-deterministic behavior of the tested LLMs, although some variation remains visible. The model cut-off date for all LLMs was set to June 2023 and a seed parameter of 128 was used in this and consequent experiments to make comparing the results more coherent and reproducible.
The reproduction uses the same screening task set of 20 papers from the prior SR  (Kuutila et al., 2020), as in the experiment done with humans. We constructed a prompt template following the screening procedure of the prior SR as closely as possible. Using this template, we generated prompts using the ZS prompting technique (see Section 2.5) to use in the reproduction. These prompts were then processed with a Python script, which called the OpenAI API. The prompt template, the generated prompts, the Python script, and the API responses are provided in the replication package (see Data Availability).
@@ 3.5. LLM Reproduction With Prompt Optimization
We selected 20 additional papers from the prior SR (Kuutila et al., 2020) as “seed papers”, of which 10 should be included and 10 should be excluded. The seed papers were used as examples with OS, FS, and FS-CoT prompting (see Section 2.5).
@@
For FS-CoT prompting, we reused the FS prompts. The phrase “Let’s think step by step” was appended to the end of each prompt. This phrase is used in Chain-of-Thought (CoT) prompting and improves task accuracy (Wei et al., 2023; Kojima et al., 2022).
@@ 3.8. Reproducing a Larger Screening Procedure
We also conducted exploratory analysis on 1,361 Scopus Papers that were initially gathered for the prior SR  (Kuutila et al., 2020) to see how LLMs perform with a more realistic sample size (see Figure 1). The Scopus Papers do not encompass all the screened articles for the prior SR, as the screened articles also included those from Google Scholar and articles resulting from snowballing, but it is the data that had complete abstracts readily available for analysis. Articles without abstracts or to which we did not have access were removed, leaving us with 1,306 papers.
@@ 4.1.2. Exploratory Analysis
Table 7. Human screener majority vote per paper. Paper references are given in the replication package  (Huotala et al., 2024).
@@ 4.2. LLM Screening
Correct (%) Model Exclude Include Precision Recall
@@
Reference: Human performance
@@ 5. Discussion
Paid researchers outperformed students in the screening task. This result is partly in conflict with the results of the experiment conducted by Salman et al. (Salman et al., 2015), where it was concluded that student performance was similar to professional performance in a programming task when the approach used was novel for both groups. In our case, the students had been taught about conducting secondary studies during the course they were attending, thus it was not completely novel for them, we did not record the experience of the researchers in performing them. Nevertheless, a noticeable result from the experiment with human screeners is that being a researcher and having higher scientific literacy skills, measured with the TOSLS, has a predictive ability on the number of correct judgments according to our regression model.
@@
The two existing ML tools, Abstrackr and Research Screener (covered in Section 2.4) had a screening accuracy between 16% and 45% (Chai et al., 2021; Gates et al., 2018; Rathbone et al., 2015). Compared to our exploratory testing on a larger screening procedure, in Table 9, our screening performance seems significantly higher than with Abstrackr and Research Screener. However, the results are not directly comparable, as the datasets used were different, and Abstrackr gives recommendations instead of judgments. Also of note is that we picked 20 screening tasks for our experiments, and the performance for these particularly difficult cases is between 50% and 75% with a recall between 0.3 to 0.9. Precision and recall for each LLM and prompt combination is given in the replication package (Huotala et al., 2024). The recent studies on LLMs for automating the screening process of SRs (Guo et al., 2024; Khraisha et al., 2024; Bolanos et al., 2024; Wilkins, 2023; Robinson et al., 2023; Wang et al., 2024) show that the accuracy of title-abstract screening with LLMs have varied between 0.60 and 0.835, which are similar to our results. Similar to the studies done with Abstrackr and Research Screener, our results are not directly comparable with the recent studies on LLMs, because of the datasets that focus on different aspects. Additionally, not all studies mention the used LLM versions, seed or temperature parameters making comparisons harder.
@@ 6. Limitations
As conducting SRs includes more than just paper screening, another avenue for future work is considering the use of LLMs in more phases. This includes the identification of studies for the screening (Wohlin et al., 2022), data extraction and analysis.
@@ 7. Conclusions
The datasets and source code for the experiments are available at Zenodo  (Huotala et al., 2024). However, the data involving human screeners is not shared due to the absence of explicit consent for distribution.
Acknowledgements.